Calculate the molecular masses of   CO2
enweuvbehwvbfyehwbuvjrnve

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


CO2⇒ 1 × 12 + 2 × 16 = 12 +32 = 44 u



Thanks